# literature.bib

### v1

@@ 8 OPDI_BARRIER()
Listing 3.1: OpenMP parallel code written according to SU2’s macro interface (top left) and the corresponding final macro expansion (top right). Hybrid AD compatible code written according to the extended macro interface (bottom left) and the corresponding intermediate macro expansion to OpDiLib macros (bottom right).
@@
CoDiPack has already been applied to SU2 for the purpose of serial and, together with MeDiPack, MPI-parallel discrete adjoints [AlbringSG2016]. SU2 already supports the exchange of the floating point computation type by means of a global typedef su2double. An AD tool wrapper, implemented by functions in a namespace AD, abstracts the interaction with CoDiPack. Furthermore, an OpenMP abstraction layer provides SU2_OMP_* macros that replace OpenMP #pragma directives throughout the SU2 code, an example is given in the top part of LABEL:listing:su2_macro_expansion_opdi. We build on these facilities for enabling hybrid parallel discrete adjoints.
We update SU2 to use version 2.1 of CoDiPack, which introduces stable support for OpDiLib. Following the OpenMP parallelism in $G$ and $J$, OpDiLib allows us to record their computational graphs in an OpenMP parallel fashion and employ a corresponding parallelism for the evaluation of the derivative. As detailed in [BluehdornSG2023], OpDiLib features two strategies (backends) for the detection of OpenMP constructs so that they can be augmented and differentiated accordingly: a fully automatic approach that leverages the OpenMP Tools Interface (OMPT) [EichenbergerEtAl2013], and a semi-automatic approach that, in particular, requires using OPDI_* macros instead of OpenMP #pragma directives, while preserving compatibility with the OMPT backend for compilers that support OMPT. We pursue the semi-automatic approach so that hybrid parallel discrete adjoints in SU2 can make use of OMPT but do not depend on it. We revise the SU2_OMP_* macros so that they admit expansion to a corresponding OPDI_* macro. To match OpDiLib’s OPDI_NOWAIT replacement clause for nowait, we introduce SU2_NOWAIT. As OpDiLib requires corresponding OPDI_END_* macros after each OpenMP directive that is followed by a structured block, we also introduce corresponding END_SU2_OMP_* macros that expand to the OPDI_END_* macros and apply them throughout the SU2 code. The bottom part of LABEL:listing:su2_macro_expansion_opdi displays an example for hybrid AD compatible OpenMP parallel code in SU2 and the corresponding intermediate expansion to OpDiLib macros. When building SU2 with hybrid AD support, the OpDiLib backend is chosen automatically according to the detected compiler support for OMPT. Optionally, the user can override this with the opdi-backend build option.
We perform the initialization and finalization of OpDiLib in new omp_initialize() and omp_finalize() functions in the OpenMP abstraction layer. The AD tool wrapper becomes aware of OpDiLib’s internal states, which complement tapes and evolve alongside the recordings. We match exports and recoveries of tape positions by exports and a recoveries of OpDiLib’s state, respectively. We account for OpDiLib-specific preparations and cleanups by notifying OpDiLib prior to and after tape evaluations. As tapes are thread-local objects that change at runtime, we resolve all tape references dynamically via AD::getTape().
@@
For serial and MPI-parallel discrete adjoints, SU2 defines su2double to be CoDiPack’s type codi::RealReverse. For hybrid parallel discrete adjoints, we exchange this with the OpenMP-compatible CoDiPack type codi::RealReverseIndexOpenMP (more precisely, an instance of the template codi::RealReverseIndexOpenMPGen that ensures non-atomic evaluations for serial parts of the code, see Section 4). An important difference between these types is the underlying strategy for managing identifiers. codi::RealReverse features a linear management scheme, where identifiers are assigned in an incremental fashion and each identifier is assigned at most once throughout a recording. codi::RealReverseIndexOpenMPGen relies on an OpenMP-aware reuse management scheme, where each thread maintains a pool of identifiers from which new identifiers are drawn and to which identifiers of destructed or overwritten variables are released. Identifiers can be associated with different variables throughout a recording as long as an identifier is not associated with multiple variables at the same time. An overview of identifier management schemes and their properties is given in [SagebaumBG2020]. We identify and eliminate code patterns in SU2 that are not compatible with reuse management schemes, making the discrete adjoint in SU2 at the same time less dependent on specific AD tool properties. We introduce a codi-tape build option that can be used to test compatibility with reuse management schemes also in a non-hybrid setting.
@@
We remark that the taping strategy for SU2’s multiphysics solvers [BurghardtEtAl2021] is not compatible with reuse management schemes yet in the case that more than one zone is used. This is due to the joint recording of contributions from multiple zones on a single tape in an interleaved manner and a partial tape evaluation strategy that may skip parts of the tape. Reused identifiers can be associated with different variables in different parts of the tape, which is reflected in the tape evaluation by setting left hand side adjoint variables to zero, like the $\bar{w}=0$ in the context of (2.6). If an evaluation omits parts of the tape, it also omits the corresponding adjoint variable resets, so that contributions to different adjoint variables are mixed at the same memory location.
@@
As outlined in Section 2, $G$ relies on subroutines to solve LSEs, which are differentiated analytically by means of solving an LSE with the transposed matrix. Previously, SU2 provided dedicated routines for the multiplication with the transposed matrix, without actually transposing the matrix. This requires specialization with respect to each preconditioner and is inefficient to parallelize with OpenMP as it involves strided memory access. Therefore, we transpose the matrix in place instead, so that the primal LSE solver routines can be reused, including their parallelism and all preconditioners. The improved memory access pattern and increased flexibility should outweigh the cost of transposing the matrix.
@@
In hybrid AD builds, we disable the preaccumulations for which we cannot guarantee no-sharing. We do this either locally, or, for larger parts of the code, in particular edge loops that are parallelized according to the reduction strategy (see Section 4), by means of AD::PausePreaccumulation() and AD::ResumePreaccumulation() functions, which we introduce in SU2’s AD tool wrapper.
@@
Discrete adjoint driver, iteration, and solver classes implement steps of the reverse accumulation AD workflow summarized in Section 2. Tape recordings and evaluations in a discrete adjoint solver inherit the underlying primal solver’s OpenMP parallelism. In addition, we parallelize the (re)computation of dependent quantities after AD input registration as well as the reverse accumulation AD workflow itself. We introduce parallel loops for resetting input identifiers and for copying derivatives, in particular the extraction and initialization of adjoints in between iterations. This requires lookups in the vector of adjoint variables according to previously remembered identifiers of input and output variables. We revise the storage strategy for these identifiers. Previously, identifiers of input and output variables were stored upon AD input registration and output registration, respectively, in a sequential manner in large arrays, and later retrieved with the assumption that the order of variable accesses is always the same. Instead, we store identifiers alongside solution variables in accordance with the revised storage layout [GomesEP2021]. This does not only clarify the relation between identifiers and variables but is also compatible with OpenMP parallel extraction and recovery of identifiers.
@@
We were able to hide many of the changes discussed in the previous paragraphs in SU2’s abstraction layers for AD and OpenMP, which is in line with the principle that general SU2 development should not require in-depth familiarity with these topics. There are a two notable exceptions. First, the END_SU2_* macros throughout the SU2 code are uncommon for OpenMP development in C++. To mitigate their maintenance effort, we introduce automatic syntax checks as part of SU2’s hybrid AD build tests. These checks detect both missing end macros as well as mismatched pairs of opening and closing macros. Second, working with preaccumulations and the shared reading optimization in the context of hybrid AD requires certain awareness of the data access patterns. Declared no-sharing, if violated, can become a new source of data races as the SU2 code is developed further.
Data races can be difficult to identify and debug, in particular if they only occur during tape evaluations, where the relation to the underlying error in the recorded code is often unclear. We therefore apply the thread sanitizer, a tool for the automatic detection of data races [SerebryanyI2009]. We conduct a comprehensive analysis of SU2 and fix various data races both in the primal and in the discrete adjoint code, including violations of declared no-sharing. In the course of this, we introduce SU2_OMP_SAFE_GLOBAL_ACCESS macros that abbreviate frequently occurring barrier-master-barrier patterns within parallel regions.
To detect data races as early as possible and to mitigate the maintenance effort of hybrid parallel code, both with and without AD, we extend SU2’s CI pipeline to automate the thread sanitizer analysis. We provide thread-sanitizer enabled containers as part of the su2code/Docker-Builds repository [^5] https://github.com/su2code/Docker-Builds, which are not only the basis for the CI pipeline tests but can also be used to test for data races locally, with notably reduced setup and configuration effort.
@@ 4 Performance Optimizations
This affects in particular the management parts of SU2’s AD workflow, like the seeding and extraction of adjoint variables between discrete adjoint iterations (2.2). Additionally, we improve the preaccumulation performance by implementing the same improvements in CoDiPack for local preaccumulation workflows.
@@
Consider two statements in the primal code in the spirit of (2.5) that are executed and recorded in parallel and have a common variable on the right hand side,
@@
The corresponding adjoint updates on $\bar{u}$ according to (2.6) read
@@
As the primal parallelism is mirrored in the tape evaluation, the adjoint updates are also executed in parallel, which leads to a data race on $\bar{u}$. Atomic increments on adjoint variables resolve these data races, see, e. g., [Foerster2014, FoersterNU2011]. Parts of the parallel code where different threads do not share variables on right hand sides (referred to as exclusive read property in [Foerster2014]), on the other hand, do not require atomic updates. To leverage developer knowledge about the data access patterns, we introduce new AD::StartNoSharedReading() and AD::EndNoSharedReading() functions in SU2’s AD tool wrapper and use them to mark such parts of the code. Internally, they change OpDiLib’s adjoint access mode to use non-atomic increments, and thus avoid the overhead of atomic updates. We remark that neither serial parts of the recording (anything outside parallel regions) nor preaccumulations use atomics as both follow CoDiPack’s evaluation procedure associated with the AD type (as opposed to OpDiLib’s evaluation procedure for parallel regions that can switch between atomic and non-atomic updates). This is valid for preaccumulations due to their strict no-sharing requirement. In addition, we declare serial parts within parallel regions that are isolated in a barrier-master-barrier fashion as safe for non-atomic evaluation. The shared reading optimization is enabled by default. We add an opdi-shared-read-opt build option to disable it, e. g., to study its impact on the evaluation performance.
@@
Further potential to eliminate atomic updates lies in SU2’s strategies for loop parallelization. Given the discretization (i. e. mesh) of a computational domain, implementations of the finite volume or finite element methods require loops over the discretization elements (control volumes, elements, edges, points), to apply the numerical schemes used to solve the discrete problem. The two main aspects of enabling shared-memory parallelism of such loops are adding the required work-sharing directives to the code, and developing strategies to avoid data races that would otherwise result from gather-scatter access patterns. For example, in finite-volume solvers, the flux across a face (or edge) depends on the quantities of two adjacent control volumes, and in turn, this flux contributes to the divergence (also known as residual) of both control volumes. Of the strategies evaluated in [GomesEP2021], two were chosen for loops with high arithmetic intensity, namely grid coloring and reductions. In the former, edges and elements are colored such that data races do not occur when threads loop simultaneously over entities with the same color (synchronization is required once per color). The reduction strategy is a type of scatter-to-gather loop transformation that consists in first storing the results of a loop over edges (without scattering) and then reducing this intermediate variable by looping over the faces of each control volume. Coloring reduces the locality of the data, whereas the reduction strategy increases storage requirements and introduces a relatively expensive reduction (approximately the cost of one linear solver iteration) without sacrificing locality or parallelism. The caching performance of coloring is improved by using color groups. Within a group, the no-sharing constraint is lifted, so that all groups of a color can be processed in parallel, but each group must be assigned to a single thread. As a rule of thumb, larger color groups result in better caching, but make it more difficult to obtain a coloring with well-balanced parallel workloads.
@@
With a view on the primal performance, SU2 switches to the reduction strategy on all MPI ranks on which coloring with the specified edge color group size fails or yields an efficiency below 0.875 in terms of load balancing across the threads of the rank [GomesEP2021]. The tradeoffs between coloring and reductions are different for discrete adjoints. As was also observed in [HueckelheimHSM2019] in the context of edge loop parallelization in a flow solver, the coloring approach leads to exclusive read access. We can therefore apply the shared reading optimization. The reduction strategy, on the other hand, does not only have inherent simultaneous read accesses and cannot be optimized in this regard but also results in larger tapes that consume more memory and take longer to evaluate. For these reasons, coloring is preferential for discrete adjoints even at the cost of reducing the edge color group size. To increase the success rate of the coloring algorithm, we allow larger numbers of colors by default. To find a coloring that both performs well in terms of caching and is efficient in terms of load balancing, each MPI rank adapts its edge color group size. We assume a monotonic relationship between edge color group size and efficiency and treat coloring success and coloring efficiency of at least 0.875 as constraints. Starting with the specified edge color group size as an upper bound, each MPI ranks determines the largest admissible edge color group size by means of bisections. This adaptive strategy should result in better caching than resorting to coloring without groups (edge color group size one) and is at the same time more flexible and automatic than manual tuning of the edge color group size. The decisions are supported by the performance study in Section 5.2. Adaptivity is enabled by default and can be toggled with the EDGE_COLORING_RELAX_DISC_ADJ SU2 configuration file option.
5 Benchmarking

### v2

@@ 8 OPDI_BARRIER()
Listing 2: OpenMP parallel code written according to SU2’s macro interface (top left) and the corresponding final macro expansion (top right). Hybrid AD compatible code written according to the extended macro interface (bottom left) and the corresponding intermediate macro expansion to OpDiLib macros (bottom right).
@@
CoDiPack has already been applied to SU2 for the purpose of serial and, together with MeDiPack, MPI-parallel discrete adjoints [25]. SU2 already supports the exchange of the floating point computation type by means of a global typedef su2double. An AD tool wrapper, implemented by functions in a namespace AD, abstracts the interaction with CoDiPack. Furthermore, an OpenMP abstraction layer provides SU2_OMP_* macros that replace OpenMP #pragma directives throughout the SU2 code, an example is given in the top part of LABEL:listing:su2_macro_expansion_opdi. We build on these facilities for enabling hybrid parallel discrete adjoints.
We update SU2 to use version 2.1 of CoDiPack, which introduces stable support for OpDiLib. Following the OpenMP parallelism in $G$ and $J$, OpDiLib allows us to record their computational graphs in an OpenMP parallel fashion and employ a corresponding parallelism for the evaluation of the derivative. As detailed in [31], OpDiLib features two strategies (backends) for the detection of OpenMP constructs so that they can be augmented and differentiated accordingly: a fully automatic approach that leverages the OpenMP Tools Interface (OMPT) [37], and a semi-automatic approach that, in particular, requires using OPDI_* macros instead of OpenMP #pragma directives, while preserving compatibility with the OMPT backend for compilers that support OMPT. We pursue the semi-automatic approach so that hybrid parallel discrete adjoints in SU2 can make use of OMPT but do not depend on it. We revise the SU2_OMP_* macros so that they admit expansion to a corresponding OPDI_* macro. To match OpDiLib’s OPDI_NOWAIT replacement clause for nowait, we introduce SU2_NOWAIT. As OpDiLib requires corresponding OPDI_END_* macros after each OpenMP directive that is followed by a structured block, we also introduce corresponding END_SU2_OMP_* macros that expand to the OPDI_END_* macros and apply them throughout the SU2 code. The bottom part of LABEL:listing:su2_macro_expansion_opdi displays an example for hybrid AD compatible OpenMP parallel code in SU2 and the corresponding intermediate expansion to OpDiLib macros. When building SU2 with hybrid AD support, the OpDiLib backend is chosen automatically according to the detected compiler support for OMPT. Optionally, the user can override this with the opdi-backend build option.
We perform the initialization and finalization of OpDiLib in the new omp_initialize() and omp_finalize() functions in the OpenMP abstraction layer. The AD tool wrapper becomes aware of OpDiLib’s internal states, which complement tapes and evolve alongside the recordings. We match exports and recoveries of tape positions by exports and a recoveries of OpDiLib’s state, respectively. We account for OpDiLib-specific preparations and cleanups by notifying OpDiLib prior to and after tape evaluations. As tapes are thread-local objects that change at runtime, we resolve all tape references dynamically via AD::getTape().
@@
For serial and MPI-parallel discrete adjoints, SU2 defines su2double to be CoDiPack’s type codi::RealReverse. For hybrid parallel discrete adjoints, we rely on the OpenMP-compatible CoDiPack type codi::RealReverseIndexOpenMP (more precisely, an instance of the template codi::RealReverseIndexOpenMPGen that ensures non-atomic evaluations for serial parts of the code, see Section 3.2). An important difference between these types is the underlying strategy for managing identifiers. codi::RealReverse features a linear management scheme, where identifiers are assigned in an incremental fashion and each identifier is assigned at most once throughout a recording. codi::RealReverseIndexOpenMPGen relies on an OpenMP-aware reuse management scheme, where each thread maintains a pool of identifiers from which new identifiers are drawn and to which identifiers of destructed or overwritten variables are released. Identifiers can be associated with different variables throughout a recording as long as an identifier is not associated with multiple variables at the same time. An overview of identifier management schemes and their properties is given in [34]. We identify and eliminate code patterns in SU2 that are not compatible with reuse management schemes, making the discrete adjoint in SU2 at the same time less dependent on specific AD tool properties. We introduce a codi-tape build option that can be used to test compatibility with reuse management schemes also in a non-hybrid setting.
@@
We remark that the taping strategy for SU2’s multiphysics solvers [23] is not compatible with reuse management schemes yet in the case that more than one zone is used. This is due to the joint recording of contributions from multiple zones on a single tape in an interleaved manner and a partial tape evaluation strategy that may skip parts of the tape. Reused identifiers can be associated with different variables in different parts of the tape, which is reflected in the tape evaluation by setting consumed adjoint variables to zero, like the $\bar{w}=0$ in the context of (6). If an evaluation omits parts of the tape, it also omits the corresponding adjoint variable resets, so that contributions to different adjoint variables are mixed at the same memory location.
@@
As outlined in Section 2.2, $G$ relies on subroutines to solve LSEs, which are differentiated analytically by means of solving an LSE with the transposed matrix. Previously, SU2 provided dedicated routines for the multiplication with the transposed matrix, without actually transposing the matrix. This requires specialization with respect to each preconditioner and is inefficient to parallelize with OpenMP as it involves strided memory access. Therefore, we transpose the matrix in place instead, so that the primal LSE solver routines can be reused, including their parallelism and all preconditioners. The improved memory access pattern and increased flexibility should outweigh the cost of transposing the matrix.
@@
In hybrid AD builds, we disable the preaccumulations for which we cannot guarantee no-sharing. We do this either locally, or, for larger parts of the code, in particular edge loops that are parallelized according to the reduction strategy (see Section 3.2), by means of AD::PausePreaccumulation() and AD::ResumePreaccumulation() functions, which we introduce in SU2’s AD tool wrapper.
@@
In follow-up research in [38], we study the issue of simultaneous preaccumulations with shared inputs in detail, discuss approaches to re-enable such preaccumulations, and test these approaches in the context of hybrid parallel discrete adjoint computations in SU2.
@@
Discrete adjoint driver, iteration, and solver classes implement steps of the reverse accumulation AD workflow summarized in Section 2.2. Tape recordings and evaluations in a discrete adjoint solver inherit the underlying primal solver’s OpenMP parallelism. In addition, we parallelize the (re)computation of dependent quantities after AD input registration as well as the reverse accumulation AD workflow itself. We introduce parallel loops for resetting input identifiers and for copying derivatives, in particular the extraction and initialization of adjoints in between iterations. This requires lookups in the vector of adjoint variables according to previously remembered identifiers of input and output variables. We revise the storage strategy for these identifiers. Previously, identifiers of input and output variables were stored upon AD input registration and output registration, respectively, in a sequential manner in large arrays, and later retrieved with the assumption that the order of variable accesses is always the same. Instead, we store identifiers alongside solution variables in accordance with the revised storage layout [30]. This does not only clarify the relation between identifiers and variables but is also compatible with OpenMP parallel extraction and recovery of identifiers.
@@
We were able to hide many of the changes discussed in the previous paragraphs in SU2’s abstraction layers for AD and OpenMP, which is in line with the principle that general SU2 development should not require in-depth familiarity with these topics. There are a two notable exceptions. First, the END_SU2_* macros throughout the SU2 code are uncommon for OpenMP development in C++. To mitigate their maintenance effort, we introduce automatic syntax checks as part of SU2’s hybrid AD build tests. These checks detect both missing end macros as well as mismatched pairs of opening and closing macros. Second, treating preaccumulations in the presence of sharing as well as applying the shared reading optimization discussed in Section 3.2 require certain awareness of the data access patterns. Assumed or declared no-sharing, if violated, can become a new source of data races as the SU2 code is developed further.
Data races can be difficult to identify and debug, in particular if they only occur during tape evaluations, where the relation to the underlying error in the recorded code is often unclear. We therefore apply the thread sanitizer, a tool for the automatic detection of data races [39]. We conduct a comprehensive analysis of SU2 and fix various data races both in the primal and in the discrete adjoint code, including violations of declared no-sharing. In the course of this, we introduce SU2_OMP_SAFE_GLOBAL_ACCESS macros that abbreviate frequently occurring barrier-master-barrier patterns within parallel regions.
To detect data races as early as possible and to mitigate the maintenance effort of hybrid parallel code, both with and without AD, we extend SU2’s CI pipeline to automate the thread sanitizer analysis. We provide thread-sanitizer enabled containers as part of the su2code/Docker-Builds repository [^5] https://github.com/su2code/Docker-Builds (visited on July 16th, 2024), which are not only the basis for the CI pipeline tests but can also be used to test for data races locally, with notably reduced setup and configuration effort.
@@ 3.2 Performance optimizations
This affects in particular the management parts of SU2’s AD workflow, like the seeding and extraction of adjoint variables between discrete adjoint iterations (2). Additionally, we improve the preaccumulation performance by implementing the same improvements in CoDiPack for local preaccumulation workflows.
@@
Consider two statements in the primal code in the spirit of (5) that are executed and recorded in parallel and have a common variable on the right hand side,
@@
The corresponding adjoint updates on $\bar{u}$ according to (6) read
@@
As the primal parallelism is mirrored in the tape evaluation, the adjoint updates are also executed in parallel, which leads to a data race on $\bar{u}$. Atomic increments on adjoint variables resolve these data races, see, e. g., [40, 41]. Parts of the parallel code where different threads do not share variables on right hand sides (referred to as exclusive read property in [40]), on the other hand, do not require atomic updates. To leverage developer knowledge about the data access patterns, we introduce new AD::StartNoSharedReading() and AD::EndNoSharedReading() functions in SU2’s AD tool wrapper and use them to mark such parts of the code. Internally, they change OpDiLib’s adjoint access mode to use non-atomic increments, and thus avoid the overhead of atomic updates. We remark that neither serial parts of the recording (anything outside parallel regions) nor preaccumulations use atomics as both follow CoDiPack’s evaluation procedure associated with the AD type (as opposed to OpDiLib’s evaluation procedure for parallel regions that can switch between atomic and non-atomic updates). This is valid for preaccumulations due to their strict no-sharing requirement. In addition, we declare serial parts within parallel regions that are isolated in a barrier-master-barrier fashion as safe for non-atomic evaluation. The shared reading optimization is enabled by default. We add an opdi-shared-read-opt build option to disable it, e. g., to study its impact on the evaluation performance.
@@
Further potential to eliminate atomic updates lies in SU2’s strategies for loop parallelization. Given the discretization (i. e. mesh) of a computational domain, implementations of the finite volume or finite element methods require loops over the discretization elements (control volumes, elements, edges, points), to apply the numerical schemes used to solve the discrete problem. The two main aspects of enabling shared-memory parallelism of such loops are adding the required work-sharing directives to the code, and developing strategies to avoid data races that would otherwise result from gather-scatter access patterns. For example, in finite-volume solvers, the flux across a face (or edge) depends on the quantities of two adjacent control volumes, and in turn, this flux contributes to the divergence (also known as residual) of both control volumes. Of the strategies evaluated in [30], two were chosen for loops with high arithmetic intensity, namely grid coloring and reductions. In the former, edges and elements are colored such that data races do not occur when threads loop simultaneously over entities with the same color (synchronization is required once per color). The reduction strategy is a type of scatter-to-gather loop transformation that consists in first storing the results of a loop over edges (without scattering) and then reducing this intermediate variable by looping over the faces of each control volume. Coloring reduces the locality of the data, whereas the reduction strategy increases storage requirements and introduces a relatively expensive reduction (approximately the cost of one linear solver iteration) without sacrificing locality or parallelism. The caching performance of coloring is improved by using color groups. Within a group, the no-sharing constraint is lifted, so that all groups of a color can be processed in parallel, but each group must be assigned to a single thread. As a rule of thumb, larger color groups result in better caching, but make it more difficult to obtain a coloring with well-balanced parallel workloads.
@@
With a view on the primal performance, SU2 switches to the reduction strategy on all MPI ranks on which coloring with the specified edge color group size fails or yields an efficiency below 0.875 in terms of load balancing across the threads of the rank [30]. The tradeoffs between coloring and reductions are different for discrete adjoints. As was also observed in [18] in the context of edge loop parallelization in a flow solver, the coloring approach leads to exclusive read access. We can therefore apply the shared reading optimization. The reduction strategy, on the other hand, does not only have inherent simultaneous read accesses and cannot be optimized in this regard but also results in larger tapes that consume more memory and take longer to evaluate. For these reasons, coloring is preferential for discrete adjoints even at the cost of reducing the edge color group size. To increase the success rate of the coloring algorithm, we allow larger numbers of colors by default. To find a coloring that both performs well in terms of caching and is efficient in terms of load balancing, each MPI rank adapts its edge color group size. We assume a monotonic relationship between edge color group size and efficiency and treat coloring success and coloring efficiency of at least 0.875 as constraints. Starting with the specified edge color group size as an upper bound, each MPI ranks determines the largest admissible edge color group size by means of bisections. This adaptive strategy should result in better caching than resorting to coloring without groups (edge color group size one) and is at the same time more flexible and automatic than manual tuning of the edge color group size. The decisions are supported by the performance study in Section 4.2. Adaptivity is enabled by default and can be toggled with the EDGE_COLORING_RELAX_DISC_ADJ SU2 configuration file option.
4 Benchmarking results and discussion